Clara's boat goes 14 miles per hour. Find the speed of the current in the river if she can go 8 miles downstream in the same time as she can go 6 miles upstream.
Answer by macston(5194)

C=speed of current; T=time ; D=distance
.
Upstream:
Distance=rate x time
6 miles=(14 mph-C) x T
6/(14-C)=T
.
Downstream:
8 miles=(14mph+C) x T
8/(14+C)=T
.
6/(14-C)= 8/(14+C)
6(14+C)=8(14-C)
84+6C=112-8C
14C=28
C=2
.
ANSWER: The speed of the current is 2 mph.
.
CHECK:
6/(14-C)=8/(14+C)
6/(14-2)=8/(14+2)
6/12=8/16
1/2=1/2
.